# IndiBias: A Benchmark Dataset to Measure Social Biases in Language Models for Indian Context

### v1

@@ Appendix C Prompts Used
Appendix D Resources for Intersectional SEAT measurement
@@
Appendix E Challenges in Adapting CrowS-Pairs to Indian version of CrowS-Pairs
@@
Appendix F Tuple Dataset for Positive and Negative Stereotypes in Indian Society
As outlined in Section 4.2: Bias Tuple Creation (4.2) of the main paper, we employed ChatGPT to generate tuples in the format (identity term, attribute). These tuples were then assessed by three annotators to determine whether they represented common stereotypical associations within Indian society. Table 7 showcases examples of the selected tuples, accompanied by the number of annotators who identified them as stereotypical, as indicated in the Score column. These filtered tuples can be employed to construct sentences both in English and Hindi using templates.
Appendix G Other Experiments
Confidence score for a given pair of sentence is calculated using $score(S_{1})/score(S_{2})$. A confidence score of 1 denotes that the model is unbiased. Figure 5 and figure 6 shows the KDE plot distribution obtained for confidence scores using English and Hindi sentences respectively. The multilingual bert model has the highest density around the confidence score of 1 in both cases. The models exhibit a broader range of confidence scores for Hindi sentences compared to English sentences.
Figure 5: KDE-Plot of Confidence scores for English Sentences
Figure 6: KDE-Plot of Confidence scores for Hindi Sentences
@@ Appendix G Other Experiments
Test ($\downarrow$) / Model ($\rightarrow$) XLMR IndicBert Muril Bernice mT5 mBART Bloom mGPT Llama-v2 Mistral XLMR IndicBert Muril Bernice mT5 mBART Bloom mGPT Llama-v2 Mistral
@@
Table 8: Intersectional SEAT scores (Effect sizes) for Gender-Age axis. P/U: Pleasant/Unpleasant. The underlined values indicate significant at $p=0.01$.

### v2

@@ Appendix F Prompts Used
Appendix G Resources for Intersectional SEAT measurement
@@
Appendix H Challenges in Adapting CrowS-Pairs to Indian version of CrowS-Pairs
@@
Appendix I Tuple Dataset for Positive and Negative Stereotypes in Indian Society
As outlined in Section 4.2: Bias Tuple Creation (4.1) of the main paper, we employed ChatGPT/InstructGPT to generate tuples in the format (identity term, attribute). These tuples were then assessed by three annotators to determine whether they represented common stereotypical associations within Indian society. Table 8 showcases examples of the selected tuples, accompanied by the number of annotators who identified them as stereotypical, as indicated in the Score column. These filtered tuples were employed to construct sentences both in English and Hindi using templates as described in section 4.2.2. We have only generated Crows-Pairs style sentences corresponding to the Religion and Caste category using these curated tuples.
@@ Appendix I Tuple Dataset for Positive and Negative Stereotypes in Indian Society
Test ($\downarrow$) / Model ($\rightarrow$) XLMR IndicBert Muril Bernice mT5 mBART Bloom mGPT Llama-v2 Mistral XLMR IndicBert Muril Bernice mT5 mBART Bloom mGPT
@@
Table 9: Intersectional SEAT scores (Effect sizes) for Gender-Age axis. P/U: Pleasant/Unpleasant. The underlined values indicate significant at $p=0.01$. NOTE: Hindi language data are not there in the pre-training corpus of Llama-v2 and Mistral.
@@
Figure 7: Examples of paired instances (S1-S2) from Indian Crows-pairs corpus created using tuple dataset (refer, section 4.2.2). S1 always presents a stereotype or an anti-stereotype (contradicting the stereotype) for the corresponding bias tuple. The bias tuple (identity term, stereotypical attribute) for each example pair is mentioned below the bias type within parenthesis. S2 is created with minimal edits to S1 so that the identity term in S2 corresponds to a comparable advantaged group for the attribute in focus. Pairs where S1 expresses a stereotype are called "stereo", while those expressing an anti-stereotype are called "anti-stereo". The Hindi examples mentioned here are the Hindi version of the corresponding English pair.